OpenStudy (anonymous):

The sum of two numbers is 15, and their difference is 3. What are the numbers?

OpenStudy (anonymous):

\[ a+b=15\\ a-b=3 \]

OpenStudy (anonymous):

Adding both equations gets you: \[ 2a = 18 \implies a=9 \]

OpenStudy (anonymous):

\[ 9+b=15 \implies b=6 \]

OpenStudy (anonymous):

Sure enough\[ 9+6=15\\ 9-6 = 3 \]

OpenStudy (anonymous):

you write two equations and solve for y, then substitute the answer into the second equation! x + y = 15, x-y = 3 y = 15-x x - (15-x) = 3 x - 15 + x = 3 2x - 15 = 3 x = 9 15-9 = 6

OpenStudy (anonymous):

whoops nvm LOl too late xD

OpenStudy (anonymous):

thank you all :D

OpenStudy (unklerhaukus):

\[\left[\begin{array}{cc|c}1&1&15\\1&-1&3\end{array}\right]\]\[\sim\left[\begin{array}{cc|c}1&1&15\\0&-2&-12\end{array}\right]\qquad R_2\to R_2-R_1\]\[\sim\left[\begin{array}{cc|c}1&1&15\\0&1&6\end{array}\right]\qquad R_2\to R_2/-2\]\[\sim\left[\begin{array}{cc|c}1&0&9\\0&1&6\end{array}\right]\qquad R_1\to R_1-R_2\]